OpenStudy (anonymous):

a ball is dropped from a heiht of 100 feet one second later another ball is dropped from a height of 75 feet .which ball hits the ground first???

OpenStudy (anonymous):

equation for first is \[h(t)=100-16t^2\]

OpenStudy (anonymous):

after one second it is 84 feet yes?

OpenStudy (anonymous):

and it's velocity is \[h'(1)=-32\]

OpenStudy (anonymous):

so i guess the new equation, starting at t = 1 is \[h(t)=84-32t-16t^2\]

OpenStudy (anonymous):

other one has equation \[h(t)=75-16t^2\]

OpenStudy (anonymous):

see which gets to zero first

OpenStudy (anonymous):

if i knew would i ask you???

OpenStudy (anonymous):

what i mean is set each equation above = 0 and solve.

OpenStudy (anonymous):

whatever is time is least is the first to get there. i.e. solve \[84-32t-16t^2=0\] and \[75-16t^2=0\]

OpenStudy (anonymous):

first one gives \[t=\frac{3}{2}\]

OpenStudy (anonymous):

second one gives \[\frac{5\sqrt{3}}{4}\] which is bigger than 1.5

OpenStudy (anonymous):

so first ball hits ground first\??

OpenStudy (anonymous):

1st ball falls in 4.47s(approx) 2nd in 3.8s. so if the 2nd ball is thrown 1 second later... the 1st ball should fall first.